# Enhancing the Sensitivity to Seesaw Predictions in Gauged $B-L$ Scenarios

### v2

@@ I Introduction
The observation of neutrino flavor oscillations implies the existence of neutrino masses. While a large number of possible scenarios can account for neutrino masses, all of them require physics beyond the Standard Model (SM). A particularly well-studied possibility is the so-called “seesaw” scenario in which new right-handed sterile neutrinos mass mix with the left-handed neutrinos in such a way as to explain the smallness of the neutrino masses Minkowski:1977sc ; Gell-Mann:1979vob ; Mohapatra:1979ia . The existence of such states may be further motivated by new gauge symmetries. In particular, the combination of baryon and lepton number, $B-L$, is non-anomalous in the SM, provided there exist three additional right-handed neutrinos PhysRevD.20.776 ; Mohapatra:1980qe , making it a favorable extension of the SM. Thus, gauging the combination of $B-L$ can provide yet additional motivation, beyond neutrino masses, for the necessity of right-handed sterile neutrinos.
@@
where $f=L,e_{R},Q,D,u_{R},d_{R}$ for generations $i=1,2,3$. Here $Q_{f}$ are the $B-L$ charges ($-1$ for the leptons and $+1/3$ for the quarks). Diagonalizing the mass matrix of the neutrinos gives rise to an extended PMNS mixing matrix that includes the elements $U_{e,\mu,\tau}$ between the right-handed and left-handed neutrinos of each flavor. Models of gauged $B-L$ such as these that sit below the TeV scale are possible with, for example, $SO(10)$ embedding of $B-L$ Malinsky:2005bi , or leptogenesis scenarios Abada:2018oly , and the motivation for a comprehensive search for HNLs down to the $\sim$ MeV mass scale is clear Coloma:2015pih ; Borzumati:2000mc ; Acero:2022wqg . In this work, we take a relatively model agnostic set of parameters to search for HNLs in the MeV–GeV mass range, with the relevant physical phenomena described in terms of $g_{B-L}$, $U_{\alpha}$, $m_{Z^{\prime}}$, and $m_{N}$.
@@
Ordinarily, the mixings $U_{\alpha}$ alone can give rise to the production of sterile neutrinos from any process that would also produce active neutrinos. For example, in accelerator targets, one may produce heavy sterile neutrinos from the decays of charged pions, the charged and neutral kaons, and muons, which may take place via the muon and electron mixings, $U_{\mu,i}$ and $U_{e,i}$, respectively. Tau mixings $U_{\tau,i}$, on the other hand, could give rise to sterile neutrino production from tau lepton decays and $D_{s}$ meson decays Ballett:2019bgd . The sterile neutrino or HNL decay could then proceed via the same mixings in electroweak-mediated decays.
@@
If the heavy neutrino states are connected to a broken $U(1)_{B-L}$ symmetry, the massive $Z^{\prime}$ gauge boson could, therefore, contribute additional production mechanisms for the sterile neutrinos if the gauge coupling is not too small. It is this possibility which is the focus of the present study. In this work, we investigate the contribution to enhanced HNL production due to the production of a new vector, $Z^{\prime}$, which could be produced from proton bremsstrahlung, electron/positron bremsstrahlung, electron/positron annihilation, and neutral meson decays in the proton beam target experiments. We adopt this heuristic model setup of a $Z^{\prime}$ with gauge coupling $g_{B-L}$ and a single HNL $N$ with mixings to the active neutrinos $U_{\alpha}$ and calculate the modified model parameter space sensitivity of ongoing and future fixed target experiments; namely, SBND SBND , MicroBooNE microboone , and ICARUS icarus using the booster neutrino beam (BNB), as well as the future DUNE experiment and its near detector (ND) Jena:2024ycp , and the DarkQuest experiment that enjoys a much shorter distance from the beam target to detector darkquest . Similar studies focused on probing gauged $B-L$ with HNLs at the LHC and SHiP Batell:2016zod ; Deppisch:2019kvs have indicated that sensitivity to the seesaw parameter space is possible, and in the present work we show that the aforementioned fixed target experiments will also probe the seesaw parameter space through complimentary HNL production channels.
@@ II Production channels of $Z^{\prime}$
In proton beam-dump experiments the new $B-L$ gauge boson can be produced either directly via proton bremsstrahlung or indirectly via secondary SM particles. The secondary particles we are interested in are photons and positrons. In the context of $B-L$, each photon can be substituted by a $Z^{\prime}$ when kinematics allows it. In particular, we consider photons from neutral meson decays and electron bremsstrahlung. On the other hand, positrons can lead to resonant production of $Z^{\prime}$ through on-shell annihilation with electrons. Here we do not consider $Z^{\prime}$ production through Compton scattering Celentano:2020vtu , since it has a smaller cross section relative to the other processes and does not change our conclusions in the relevant region of the parameter space. We also neglect charged meson decays, but in § V, we compare our final results with those obtained in ref.Berryman:2019dme , where such a channel is studied in detail.
@@
In order to simulate the production of all particles in the proton beam interactions, we used the GEANT4 simulation toolkit GEANT4:2002zbu ; Allison:2006ve ; Allison:2016lfl . The implemented DUNE neutrino production target is a 1.5 m long cylindrical graphite rod with 1.7 cm diameter, following the description in the LBNF beamline design Papadimitriou:2017zai . For the BNB experiments, we take a beryllium target cylinder 193 cm long and 1 cm in diameter MiniBooNE:2008hfu . We are also interested in HNL production utilizing the BNB beam dump of dimensions a 4 m wide $\times$ 4 m tall $\times$ 4.21 m long, of which the upstream most 2.64 m long portion stainless steel, followed by a 0.91 m thick concrete and finally a 0.66 m thick stainless steel layers along the beam direction MicroBooNE:2015bmn .
@@ II.1 Proton Bremsstrahlung
First, we consider the $Z^{\prime}$ bremsstrahlung of the primary beam protons, Fig. 1. The diffractive scattering process is modeled in the quasi-real approximation by considering producing the vector $Z^{\prime}$ as initial state radiation (see ref. Foroughi-Abari:2021zbm for details). We briefly summarize the calculation as follows. The differential scattering cross-section is given by
@@
where $w^{D}(z,p_{T}^{2})$ is a splitting function over the longitudinal momentum fraction $z=p_{Z^{\prime}}/p$ of the outgoing $Z^{\prime}$ momentum with respect to the initial proton momentum $p$, $p_{T}$ is the transverse momentum of the outgoing $Z^{\prime}$, and $s^{\prime}=2m_{p}(p(1-z)+m_{p})$ where $m_{p}$ is the proton mass. In the construction of the splitting function $w^{D}(z,p_{T}^{2})$ in ref. Foroughi-Abari:2021zbm , there are hadronic form factors introduced to capture timelike momentum transfer as well as the off-shell momentum in Fig. 1. The parametrization of the form factors can have a significant impact on the magnitude of the cross section (up to an order of magnitude), especially in the parameter $\Lambda_{p}\sim\mathcal{O}(m_{p})$ which acts as a cutoff scale for the off-shell behavior. In this work we take the central value in the range considered by ref. Foroughi-Abari:2021zbm for the cutoff scale, $\Lambda_{p}=1.5$ GeV.
In the integration of Eq. 2, certain cuts have to be made to ensure the applicability of the quasi-real approximation; following ref. Foroughi-Abari:2021zbm , we define
@@
Assuming that most of the protons get absorbed in the proton target, we can estimate the differential rate of $Z^{\prime}$ particles produced by proton bremsstrahlung as the number of POT times the ratio of the bremsstrahlung and total proton cross sections Berryman:2019dme ; Dev:2023zts ; [^1] During the completion of this work, other analyses have proposed updated treatments of the form factors used proton bremsstrahlung, see for example ref. Foroughi-Abari:2024xlj and refs. Gorbunov:2024vrc ; Gorbunov:2023jnx . The potential impact of these new calculations is left to a future work.
@@ II.2 Neutral Meson Decay
where ${\rm Br}(\pi^{0}\to\gamma\gamma)=0.98823$ and ${\rm Br}(\eta^{0}\to\gamma\gamma)=0.3941$ Zyla:2020zbs . For each meson in the GEANT4 sample, we perform a 2-body decay monte carlo for $M\to\gamma Z^{\prime}$ in the rest frame, then boost to the lab frame to simulate the boosted spectrum of $Z^{\prime}$.
@@ II.3 Electron and Positron Bremsstrahlung
To calculate the number of $Z^{\prime}$ bremsstrahlung off of $N_{e^{\pm}}^{ij}$ electrons/positrons the $(i,j)$-th $(E_{e^{\pm}}^{i},\theta_{e^{\pm}}^{j})$ bin (where $\theta_{e^{\pm}}^{j}$ is the angle of the electron/positron with respect to the proton beam axis) as depicted in Fig. 3 left, we integrate over the GEANT4-simulated $e^{\pm}$ fluxes, convolving them with the differential cross section $\partial^{2}\sigma/(\partial E_{Z^{\prime}}\partial\Omega_{Z^{\prime}})$ for the process $e^{\pm}N\to e^{\pm}NZ^{\prime}$. The energy loss of the electrons and positrons in the material during particle transport must also be folded into the event rate calculation. The differential number flux of $Z^{\prime}$ can be expressed as
@@
where $N_{A}$ is Avogadro’s number, $X_{0}$ is the radiation length of the electrons/positrons in the dump material, $A$ is the atomic weight, and $N^{i}_{e^{\pm}}$ is the number of $e^{\pm}$ with energy $E_{e}^{i}$ and angle $\theta_{e}^{j}$ with respect to the beam axis. $I(t,E_{i},E_{f})=\frac{\theta(E_{i}-E_{f})}{E_{i}\Gamma(4t/3)}(\ln E_{i}/E_{f})^{4t/3-1}$ is the energy loss smearing function for the electron/positron radiation length $t$ integrated up to target radiation thickness $T$ PhysRevD.34.1326 . This expression gives us the differential rate of $Z^{\prime}$ over the differential angle $d\Omega_{Z^{\prime}}=d(\cos\theta_{Z^{\prime}})d\phi_{Z^{\prime}}$ with respect to the $e^{\pm}$ direction $\theta_{e^{\pm}}^{j}$; we can then transform this angle to the lab frame in which the beam axis points along the $z$-direction via
@@ II.4 Resonant Production
The prefactors and energy loss function are the same as those defined for electron/positron bremsstrahlung in § II.3. $\sigma_{\rm res}$ is the cross section for resonant production and is given in Celentano:2020vtu ;
@@
We note here that modeling the electron/positron energy loss for bremsstrahlung and resonant production using the track-length distribution function does not account for the random walk nature of the electromagnetic showers in a beam target. Recently, in ref. Blinov:2024pza showed using the PETITE package that more explicit modeling of the electromagnetic shower in the production of feebly-coupled vector particles can yield significantly different fluxes of the produced boson due to the random walk of the parent electrons/positrons deviating in their momentum direction in addition to losing energy as they transport through material. We can heuristically account for the broadening of the $Z^{\prime}$ angular distribution that would occur this way by limiting the number of track lengths integrated over in the above flux calculations. In the case of DUNE, we use a single value for $t_{\rm max}$, regardless of the original production point of positrons in the target. This average ($t_{\rm max}$) is calculated as the radiation lengths of the target downstream of the 50% positron production point. The final result is $t_{\rm max}=3.3$. In the case of the BNB dump, considering that secondary particles are propagating through the relatively thick beam dump, we take $t_{\rm max}=5$. In each case, we find an agreement with the PETITE-derived fluxes to within 20% difference.
@@
Figure 4: Number of $Z^{\prime}$ pointed within the solid angle of the DUNE near detector as a function of $m_{Z^{\prime}}$ for the different production channels, assuming $g_{B-L}=1$ and $1.47\times 10^{22}$ POT. In the case of proton bremsstrahlung, we do not use this channel below $m_{Z^{\prime}}=250$ MeV as per Foroughi-Abari:2021zbm , but show the behavior of the event rate (dashed red) here for the interest of the reader.
@@ III HNL decays in the detector
DUNE Berryman:2019dme 579 5 $7\times 3$ $4.3\times 10^{-3}$ $1.47\times 10^{22}$
@@
where $l=e,\mu,\tau\,$. For the decay width of $Z^{\prime}\to NN$, the decay width formula is calculated for right-handed Majorana $N$ final states;
@@
while the decay to left-handed neutrinos is simply
$\Gamma(Z^{\prime}\to\nu_{L}\bar{\nu}_{L})$ $=\frac{g_{B-L}^{2}m_{Z^{\prime}}}{24\pi}$ (14)
Lastly, the decays to hadrons are accessible for $m_{Z^{\prime}}>m_{\rho}=770$ MeV. We use the $R$-ratio of hadronic to muonic final states in $e^{+}e^{-}$ annihilation to parameterize the hadronic matrix elements as per refs. Ilten:2018crw ; Ezhela:2003pp ; Bauer:2018onh ;
@@
Now we discuss the possibility of HNL decays in the detector. We consider all possible decay modes of the HNL mediated by its mixing angle $U_{\alpha}$ for $\alpha=e,\mu,\tau$. There are many modes which we do not list in their entirety here but are given in refs. Coloma:2020lgy ; Abdullahi:2023gdj ; Berryman:2017twh . Some dominant decay modes and the approximate range of HNL masses over which they are relevant ($\gtrsim 5$ branching ratio) are

### v3

@@ I Introduction
The observation of neutrino flavor oscillations implies the existence of neutrino masses. While a large number of possible scenarios can account for neutrino masses, all of them require physics beyond the Standard Model (SM). A particularly well-studied possibility is the so-called “seesaw” scenario in which new right-handed sterile neutrinos mass mix with the left-handed neutrinos in such a way as to explain the smallness of the neutrino masses [1, 2, 3]. The existence of such states may be further motivated by new gauge symmetries. In particular, the combination of baryon and lepton number, $B-L$, is non-anomalous in the SM, provided there exist three additional right-handed neutrinos [4, 5], making it a favorable extension of the SM. Thus, gauging the combination of $B-L$ can provide yet additional motivation, beyond neutrino masses, for the necessity of right-handed sterile neutrinos.
@@
where $f=L,e_{R},Q,D,u_{R},d_{R}$ for generations $i=1,2,3$. Here $Q_{f}$ are the $B-L$ charges ($-1$ for the leptons and $+1/3$ for the quarks). Diagonalizing the mass matrix of the neutrinos gives rise to an extended PMNS mixing matrix that includes the elements $U_{e,\mu,\tau}$ between the right-handed and left-handed neutrinos of each flavor. Models of gauged $B-L$ such as these that sit below the TeV scale are possible with, for example, $SO(10)$ embedding of $B-L$ [6], or leptogenesis scenarios [7], and the motivation for a comprehensive search for HNLs down to the $\sim$ MeV mass scale is clear [8, 9, 10]. In this work, we take a relatively model agnostic set of parameters to search for HNLs in the MeV–GeV mass range, with the relevant physical phenomena described in terms of $g_{B-L}$, $U_{\alpha}$, $m_{Z^{\prime}}$, and $m_{N}$.
@@
Ordinarily, the mixings $U_{\alpha}$ alone can give rise to the production of sterile neutrinos from any process that would also produce active neutrinos. For example, in accelerator targets, one may produce heavy sterile neutrinos from the decays of charged pions, the charged and neutral kaons, and muons, which may take place via the muon and electron mixings, $U_{\mu,i}$ and $U_{e,i}$, respectively. Tau mixings $U_{\tau,i}$, on the other hand, could give rise to sterile neutrino production from tau lepton decays and $D_{s}$ meson decays [11]. The sterile neutrino or HNL decay could then proceed via the same mixings in electroweak-mediated decays.
@@
If the heavy neutrino states are connected to a broken $U(1)_{B-L}$ symmetry, the massive $Z^{\prime}$ gauge boson could, therefore, contribute additional production mechanisms for the sterile neutrinos if the gauge coupling is not too small. It is this possibility which is the focus of the present study. In this work, we investigate the contribution to enhanced HNL production due to the production of a new vector, $Z^{\prime}$, which could be produced from proton bremsstrahlung, electron/positron bremsstrahlung, electron/positron annihilation, and neutral meson decays in the proton beam target experiments. We adopt this heuristic model setup of a $Z^{\prime}$ with gauge coupling $g_{B-L}$ and a single HNL $N$ with mixings to the active neutrinos $U_{\alpha}$ and calculate the modified model parameter space sensitivity of ongoing and future fixed target experiments; namely, SBND [12], MicroBooNE [13], and ICARUS [14] using the booster neutrino beam (BNB), as well as the future DUNE experiment and its near detector (ND) [15], and the DarkQuest experiment that enjoys a much shorter distance from the beam target to detector [16]. Similar studies focused on probing gauged $B-L$ with HNLs at the LHC and SHiP [17, 18] have indicated that sensitivity to the seesaw parameter space is possible, and in the present work we show that the aforementioned fixed target experiments will also probe the seesaw parameter space through complimentary HNL production channels.
@@ II Production channels of $Z^{\prime}$
In proton beam-dump experiments the new $B-L$ gauge boson can be produced either directly via proton bremsstrahlung or indirectly via secondary SM particles. The secondary particles we are interested in are photons and positrons. In the context of $B-L$, each photon can be substituted by a $Z^{\prime}$ when kinematics allows it. In particular, we consider photons from neutral meson decays and electron bremsstrahlung. On the other hand, positrons can lead to resonant production of $Z^{\prime}$ through on-shell annihilation with electrons. Here we do not consider $Z^{\prime}$ production through Compton scattering [19], since it has a smaller cross section relative to the other processes and does not change our conclusions in the relevant region of the parameter space. We also neglect charged meson decays, but in § V, we compare our final results with those obtained in ref. [20], where such a channel is studied in detail.
@@
In order to simulate the production of all particles in the proton beam interactions, we used the GEANT4 simulation toolkit [21, 22, 23]. The implemented DUNE neutrino production target is a 1.5 m long cylindrical graphite rod with 1.7 cm diameter, following the description in the LBNF beamline design [24]. For the BNB experiments, we take a beryllium target cylinder 193 cm long and 1 cm in diameter [25]. We are also interested in HNL production utilizing the BNB beam dump of dimensions a 4 m wide $\times$ 4 m tall $\times$ 4.21 m long, of which the upstream most 2.64 m long portion stainless steel, followed by a 0.91 m thick concrete and finally a 0.66 m thick stainless steel layers along the beam direction [26].
@@ II.1 Proton Bremsstrahlung
First, we consider the $Z^{\prime}$ bremsstrahlung of the primary beam protons, Fig. 1. The diffractive scattering process is modeled in the quasi-real approximation by considering producing the vector $Z^{\prime}$ as initial state radiation (see ref. [27] for details). We briefly summarize the calculation as follows. The differential scattering cross-section is given by
@@
where $w^{D}(z,p_{T}^{2})$ is a splitting function over the longitudinal momentum fraction $z=p_{Z^{\prime}}/p$ of the outgoing $Z^{\prime}$ momentum with respect to the initial proton momentum $p$, $p_{T}$ is the transverse momentum of the outgoing $Z^{\prime}$, and $s^{\prime}=2m_{p}(p(1-z)+m_{p})$ where $m_{p}$ is the proton mass. In the construction of the splitting function $w^{D}(z,p_{T}^{2})$ in ref. [27], there are hadronic form factors introduced to capture timelike momentum transfer as well as the off-shell momentum in Fig. 1. The parametrization of the form factors can have a significant impact on the magnitude of the cross section (up to an order of magnitude), especially in the parameter $\Lambda_{p}\sim\mathcal{O}(m_{p})$ which acts as a cutoff scale for the off-shell behavior. In this work we take the central value in the range considered by ref. [27] for the cutoff scale, $\Lambda_{p}=1.5$ GeV.
In the integration of Eq. 2, certain cuts have to be made to ensure the applicability of the quasi-real approximation; following ref. [27], we define
@@
Assuming that most of the protons get absorbed in the proton target, we can estimate the differential rate of $Z^{\prime}$ particles produced by proton bremsstrahlung as the number of protons on target (POT) times the ratio of the bremsstrahlung and total proton cross sections [20, 28]; [^1] During the completion of this work, other analyses have proposed updated treatments of the form factors used proton bremsstrahlung, see for example ref. [29] and refs. [30, 31]. The potential impact of these new calculations is left to a future work.
@@ II.2 Neutral Meson Decay
where ${\rm Br}(\pi^{0}\to\gamma\gamma)=0.98823$ and ${\rm Br}(\eta^{0}\to\gamma\gamma)=0.3941$ [32]. For each meson in the GEANT4 sample, we perform a 2-body decay monte carlo for $M\to\gamma Z^{\prime}$ in the rest frame, then boost to the lab frame to simulate the boosted spectrum of $Z^{\prime}$.
@@ II.3 Electron and Positron Bremsstrahlung
To calculate the number of $Z^{\prime}$ bremsstrahlung off of $N_{e^{\pm}}^{ij}$ electrons/positrons the $(i,j)$-th $(E_{e^{\pm}}^{i},\theta_{e^{\pm}}^{j})$ bin (where $\theta_{e^{\pm}}^{j}$ is the angle of the electron/positron with respect to the proton beam axis) as depicted in Fig. 3 left, we integrate over the GEANT4-simulated $e^{\pm}$ fluxes, convolving them with the differential cross section $\partial^{2}\sigma/(\partial E_{Z^{\prime}}\partial\Omega_{Z^{\prime}})$ for the process $e^{\pm}A\to e^{\pm}AZ^{\prime}$ off an atomic target $A$. The energy loss of the electrons and positrons in the material during particle transport must also be folded into the event rate calculation. The differential number flux of $Z^{\prime}$ can be expressed as
@@
where $N_{A}$ is Avogadro’s number, $X_{0}$ is the radiation length of the electrons/positrons in the dump material (g/cm2), $A$ is the atomic weight (g/mol), and $N^{i}_{e^{\pm}}$ is the number of $e^{\pm}$ with energy $E_{e}^{i}$ and angle $\theta_{e}^{j}$ with respect to the beam axis. $I(t,E_{i},E_{f})=\frac{\theta(E_{i}-E_{f})}{E_{i}\Gamma(4t/3)}(\ln E_{i}/E_{f})^{4t/3-1}$ is the energy loss smearing function, integrated over the dimensionless number of radiation lengths $t$ up to the radiation length of the target, $T$ [33]. This expression gives us the differential rate of $Z^{\prime}$ over the differential angle $d\Omega_{Z^{\prime}}=d(\cos\theta_{Z^{\prime}})d\phi_{Z^{\prime}}$ with respect to the $e^{\pm}$ direction $\theta_{e^{\pm}}^{j}$; we can then transform this angle to the lab frame in which the beam axis points along the $z$-direction via
@@ II.4 Resonant Production
The prefactors and energy loss function are the same as those defined for electron/positron bremsstrahlung in § II.3. $\sigma_{\rm res}$ is the cross section for resonant production and is given in [19];
@@
We note here that modeling the electron/positron energy loss for bremsstrahlung and resonant production using the track-length distribution function does not account for the random walk nature of the electromagnetic showers in a beam target. Recently, in ref. [34] showed using the PETITE package that more explicit modeling of the electromagnetic shower in the production of feebly-coupled vector particles can yield significantly different fluxes of the produced boson due to the random walk of the parent electrons/positrons deviating in their momentum direction in addition to losing energy as they transport through material. We can heuristically account for the broadening of the $Z^{\prime}$ angular distribution that would occur this way by limiting the number of track lengths integrated over in the above flux calculations. In the case of DUNE, we use a single value for $t_{\rm max}$, regardless of the original production point of positrons in the target. This average ($t_{\rm max}$) is calculated as the radiation lengths of the target downstream of the 50% positron production point. The final result is $t_{\rm max}=3.3$. In the case of the BNB dump, considering that secondary particles are propagating through the relatively thick beam dump, we take $t_{\rm max}=5$. In each case, we find an agreement with the PETITE-derived fluxes to within 20% difference.
@@
Figure 4: Number of $Z^{\prime}$ pointed within the solid angle of the DUNE near detector as a function of $m_{Z^{\prime}}$ for the different production channels, assuming $g_{B-L}=1$ and $1.47\times 10^{22}$ POT. In the case of proton bremsstrahlung, we do not use this channel below $m_{Z^{\prime}}=250$ MeV as per [27], but show the behavior of the event rate (dashed red) here for the interest of the reader.
@@ III HNL decays in the detector
DUNE [20] 579 5 $7\times 3$ $4.3\times 10^{-3}$ $1.47\times 10^{22}$
@@
where $l=e,\mu,\tau\,$. For the decay width of $Z^{\prime}\to NN$, the decay width formula is calculated for the heavy $N$ final states;
@@
while the decay to the light neutrinos is simply
$\Gamma(Z^{\prime}\to\nu\nu)$ $=\frac{g_{B-L}^{2}m_{Z^{\prime}}}{24\pi}$ (14)
Lastly, the decays to hadrons are accessible for $m_{Z^{\prime}}>m_{\rho}=770$ MeV. We use the $R$-ratio of hadronic to muonic final states in $e^{+}e^{-}$ annihilation to parameterize the hadronic matrix elements as per refs. [35, 36, 37];
@@
Now we discuss the possibility of HNL decays in the detector. We consider all possible decay modes of the HNL mediated by its mixing angle $U_{\alpha}$ for $\alpha=e,\mu,\tau$. There are many modes which we do not list in their entirety here but are given in refs. [38, 39, 40]. Some dominant decay modes and the approximate range of HNL masses over which they are relevant ($\gtrsim 5$ branching ratio) are
@@
The reader may also wonder if contributions to the HNL decay width can come from the $Z^{\prime}$ in addition to the mixing with electroweak diagrams. For the small gauge coupling we have adopted, $g_{B-L}\sim O(10^{-4})$, the contribution to the HNL branching ratios should go like $g_{B-L}^{2}/m_{Z^{\prime}}^{2}\ll g^{2}/m_{W}^{2}$ for most of the parameter space we have considered. Since in the following analysis we consider $m_{Z^{\prime}}\gtrsim 100$ MeV, the contribution is limited in comparison with the electroweak factor of $g^{2}/m_{W}^{2}\simeq 6.5\cdot 10^{-5}$ GeV-2.
As for the case $m_{Z^{\prime}}<m_{HNL}$, this could be another interesting possibility that we do not consider here. In this case, the production mechanism could still arise from proton bremsstrahlung to a virtual $Z^{\prime}$ that pair produces $NN$, which would be phase space suppressed relative to the on-shell $Z^{\prime}$ we consider here, or via the usual charged meson decays with the mixing $U_{\alpha}$ driving the production rate. So, the light $m_{Z^{\prime}}$ limit would likely capture the same sensitivity as the $U_{\alpha}$-only sensitivity, however with the presence of some additional HNL decay channels that now get a less suppressed contribution from the $Z^{\prime}$ in addition to the exclusively electroweak and $U_{\alpha}$-driven decay modes (e.g. $N\to Z^{\prime}\nu_{\alpha}$). These new modes may change the lifetime slightly and offer new final states for the experiments to search for. Other signals could arise from long-lived $Z^{\prime}\to\ell^{+}\ell^{-}$ decays, though these would be independent of the HNL decay signal and the HNL parameter space.